# TESSERA: Precomputed FAIR Global Pixel Embeddings for Earth Representation and Analysis

### v5

@@ A.1 Model details
$\mathcal{L}_{BT}=\sum_{i}(1-C_{ii})^{2}+\lambda_{BT}\sum_{i}\sum_{j\neq i}C_{ij}^{2}$ (7)
To further enhance the robustness of the model and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term $\mathcal{L}_{MIX}$, inspired by Bandara et al. [108]. For each training batch, this involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating a mixed view $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$. The mixing coefficient $\alpha_{mix}$ is sampled from a uniform distribution on the unit interval, $\alpha_{mix}\sim U(0,1)$. The embeddings $Z_{M}$ and $Z_{S}$ are obtained from their respective views. The mix-up loss penalizes deviations from the assumption that a linear interpolation in the input space corresponds to a linear interpolation in the embedding space.
$C_{target}^{MA}$ $=\alpha_{mix}(Z_{A})^{T}Z_{A}+(1-\alpha_{mix})(Z_{S})^{T}Z_{A}$ (8)
$C_{target}^{MS}$ $=\alpha_{mix}(Z_{A})^{T}Z_{S}+(1-\alpha_{mix})(Z_{S})^{T}Z_{S}$ (9)
$\mathcal{L}_{MIX}$ $=\frac{1}{2}(\|C^{MA}-C_{target}^{MA}\|_{F}^{2}+\|C^{MS}-C_{target}^{MS}\|_{F}^{2})$ (10)
$\mathcal{L}_{total}=\mathcal{L}_{BT}+\lambda_{mix}\mathcal{L}_{MIX}$ (11)
@@
Operationally, the process begins with the aggregation of d-pixels from all MGRS tiles into a single, comprehensive pool. A global shuffling operation is performed on this pool, a critical step to break the spatial contiguity of data from individual tiles and ensure each training batch contains a diverse mix of geographical and environmental contexts. Following this global shuffle, the data augmentation required by the Barlow Twins framework is applied. As detailed in the Supplementary Text, this involves generating two distinct augmented views (e.g., $Y^{A}$ and $Y^{B}$) for each d-pixel.
@@
Figure 10: Data shuffling greatly improves training stability. a, Schematic of the data shuffling and loading process. D-pixels (coloured squares) from thousands of MGRS tiles are first aggregated into a global pool. A custom Rust binary [37] performs a global shuffle on this multi-terabyte dataset before applying augmentations. The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches. b, Comparison of training loss curves. The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias. The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
Global Embedding Map Generation
@@
A.2 Scaling with data and network size
@@
Figure 11: Model performance improves with increasing projector size and additional computation. a, Hyperparameter sweep for the projector network’s architecture. Each point represents a configuration, with its position determined by the number of hidden layers (depth, x-axis) and their width (y-axis). The colour and size of the point indicate the validation F1 score and the number of parameters, respectively. Performance generally increases with both width and depth. The dagger (†) marks the configuration with the highest F1 score, which required a reduced batch size due to memory constraints. The star (*) indicates the final configuration selected for TESSERA v1.0, representing a balance between performance and efficiency. b, Scaling laws for TESSERA. The plot shows the validation F1 score as a function of training FLOPs and model size (in billions of parameters). A clear positive correlation is visible, indicating that performance improves as more computational resources and larger models are used. The final TESSERA v1.0 model is marked with a star (*) to demonstrate its position within this scaling trend.
@@
Figure 12: Unlike common Remote Sensing Foundation Models (RSFM), TESSERA simplifies downstream task applications. Most existing RSFMs use approach (a) where, after training, the encoder is extracted and combined with a task-specific head to form a new model for downstream tasks. Users can then fine-tune the entire model or only the head. Another approach (b) is where both the encoder and the projector (or decoder) are typically updated. In contrast, the TESSERA approach (c) uses pre-generated embeddings directly as input. This results in lightweight downstream models that consist only of a task-specific head.
Ablation study
Table 3: The ablation study demonstrates that data shuffling and mixup regularization are essential, Sentinel 1 data adds little additional value, and that quantization has nearly no impact on performance. We study downstream crop classification performance on the 100x-downsampled Austrian crop dataset. The baseline represents the full model. Each subsequent row shows the result of removing the specified component. Performance is measured by validation F1 score and the RankMe score. Arrows ($\uparrow$) indicate that higher values are better. The results demonstrate that all components contribute positively to the model’s overall performance.
Configuration Validation F1 Score ($\uparrow$) RankMe Score ($\uparrow$)
Baseline (Full Model) 0.773 0.963
w/o Global Data Shuffling 0.681 0.847
w/o Mixup Regularization 0.662 0.857
w/o Sentinel-1 Data 0.742 0.931
w/o Shuffling & Mixup 0.626 0.867
w/o Embedding Quantization 0.779 0.972
@@
The results of the ablation study are summarized in Table 3. Our analysis reveals several key insights into the model’s architecture and training strategy. The most critical component turns out to be the global data shuffling pipeline, which is designed to break spatial autocorrelation and present the model with geographically diverse mini-batches. Removing mixup regularization also severely degrades results. When both are removed, the model’s performance collapses, showing that the combination of global shuffling and mixup regularization is needed to prevent overfitting and to learn generalizable robust representations.
@@
A.3 Downstream Task Application Methodology
@@
The pipeline for applying TESSERA embeddings to downstream tasks is as follows (illustrated in the Supplementary fig. 12):
@@
    Download Embeddings for the Region of Interest: The Geotessera Python library [37] allows users to download embeddings for a desired region and year in the form of a numpy array.
@@
A.4 Downstream Task: Austrian Crop Classification
@@
We used the publicly available INVEKOS dataset for Austria, focusing on the 2022 growing season [44]. The data set originally contained 154 crop types, which we grouped into 17 broader classes (for example, combining sugar beet and feed beet) based on phenological similarity and sample availability to ensure robust training and evaluation. For each pixel in the dataset, we extracted its corresponding 128-dimensional TESSERA embedding from our generated 2022 global embeddings map.
@@
  • GSE + MLP: To include another key foundation model, we used the publicly available Google AphaEarth Satellite Embedding (GSE) V1 dataset [39]. These embeddings were similarly fed into an MLP head identical to the one used for TESSERA.
  • Random Forest: As a traditional baseline, we trained a Random Forest (RF) classifier directly on raw time-series data. For each pixel, we utilized all available Sentinel-1 (2 polarizations) and Sentinel-2 (10 spectral bands) observations throughout the year. The temporal and spectral/polarization dimensions were flattened and concatenated to form a single 1256-dimensional feature vector. The RF model was configured with 100 trees, while all other hyperparameters retained their default values as implemented in the scikit-learn library.
@@
  • For the foundation models capable of generating pixel-wise embeddings, namely TESSERA, PRESTO [19], and GSE [39], a distinct workflow was used. We first construct patches from their pixel-level embeddings (e.g., a data cube of size $H\times W\times C$). These patches were then fed into a standard UNet architecture [76], which learns to model the spatial relationships between the embeddings to produce a segmentation map. It is noteworthy that while TESSERA and PRESTO’s base embeddings do not explicitly model spatial context, GSE’s do. However, their shared ability to generate pixel-level outputs makes them suitable for this UNet-based downstream approach.
@@
Figure 13: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation. Each row corresponds to a different patch size (32x32, 64x64, 128x128). The first column shows the ground truth segmentation. For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labelled ’Embedding’) and the final predicted segmentation map (labelled ’Predicted’).
@@
The 2D visualizations shown in the main text (Figure 3 d) were generated by applying the Uniform Manifold Approximation and Projection (UMAP) [45] algorithm to the embeddings of every pixel within the Austrian study area. UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology. It constructs a weighted k-nearest neighbour graph in the original high-dimensional space, then optimizes a low-dimensional projection to primarily preserve local distances and local structure. UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space. In this work, the entire embedding map (for example, a shape array $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for the TESSERA, PRESTO, and GSE embeddings.
@@
A.5 Downstream Task: Canopy height estimation
@@
TESSERA, PRESTO [19], and GSE [39] embeddings were compared by training identical models using each representation as input, with all other settings held constant. Each model was trained for 200 epochs on the same data split, and the version achieving the best validation performance was selected for final evaluation on the held-out test set.
To benchmark performance against classical remote sensing approaches applied at the global scale, we compared the predictions against three global canopy height products: Potapov et al. [2021][59] (GLAD), Lang et al. [2023][58] (ETH), and Tolan et al. [2024][60] (Meta) as well as a regional map produced by Lang et al. [2021][68]. These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS. Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region, which is reasonable given that the area was not logged or burned.
@@
A.6 Downstream Task: Burned area detection
Study Areas and Ground Truth Datasets
For fire disturbance detection, two $\times15$ km tiles in California were selected at random from the training sample pool in the burn scars benchmark dataset [114], referred to as Burn Area A and Burn Area B. Each Burn Area contains over two million 1 m pixels. Burn Area A is a heterogeneous landscape of forests, shrublands, agricultural land, and built infrastructure. In 2020, it experienced two separate wildfires: the ’Quail Fire’ in June ($\sim$730 ha) and the much larger ’Hennessey Fire’ in August ($\sim$ 121,000 ha). Burn Area B experienced a single large fire (’Glass’) ($\sim$27,500 ha) in September 2020 that mainly affected forested land. In all cases, most of the impact of the fire occurred in vegetated areas composed of forest and shrubland. To classify the dominant land cover types in the burn areas, we used the Dynamic World Land Cover data set [115].
The dataset from the Monitoring Trends in Burn Severity (MTBS) program [95] was used to identify burned areas and severity of the burn. Fire ignition and containment dates, as well as fire perimeters, were obtained from the MTBS burned area boundaries dataset. Burn severity, defined as the measure of ecological damage caused by fire [116], was derived from categorical and continuous severity metrics provided by MTBS. To obtain burn severity categories of low, medium, and high, as well as continuous burn severity, we used the annualized MTBS burn severity dataset as well as individual severity maps for each fire. MTBS calculates severity using two primary indices: the differenced NBR (dNBR) and its relativized form (RdNBR), both calculated from pre- and post-fire Landsat images. Low severity typically indicates limited ecological impact, while high severity reflects significant ecological disruption and vegetation mortality [116]. Since MTBS datasets are derived from Landsat, they are provided with a resolution of 30 m and were upsampled to 10 m to match the resolution of the TESSERA representations. The MTBS and Dynamic World datasets were downloaded as 10 m rasters from Google Earth Engine [102]. For our classification analysis, we define “unburned” pixels using MTBS class 0 and “burned” pixels as MTBS classes 2 (low), 3 (moderate) and 4 (high). The ambiguous “unburned/low” class (1) was excluded to ensure a clear binary classification benchmark.
Dimensionality Reduction
We used UMAP (Uniform Manifold Approximation and Projection) [45] to reduce 128 dimensions of TESSERA embeddings to 2D for visual inspection of FM representations in areas affected by wildfires. We applied the dimensionality reduction to 128-dimensional TESSERA embeddings and visualized the resulting 2D projections. To interpret the structure of the projected clusters, we colour coded points according to different annotation schemes: burn state (burned vs. unburned), land cover class, or burn severity. These visualizations helped us assess whether FM embeddings spatially separate ecological conditions of interest. To reduce computational cost and improve legibility, we randomly subsampled a subset of pixels prior to applying the dimensionality reduction. UMAP projections were additionally used to classify pixels into unburned, low, moderate, and high-severity classes using a random forest (RF) classifier, to quantify the separability in the UMAP embedding space. UMAP projections were computed using the umap-learn Python library and RF was implemented using the scikit-learn Python library.
Label Efficiency Analysis and Baselines
We conducted a label efficiency experiment to evaluate the performance of a random forest classifier when only a limited amount of labelled data is available. The valid pixels from both study areas were combined and split into a fixed test set (20%) and a training set. We then trained multiple classifiers on progressively smaller stratified random subsamples of the full training set, with ratios ranging from 1.0 to 0.001%. The classifier used 100 estimators, a maximum depth of 12, minimum leaf samples of 20, and class_weight=‘balanced’ to account for class imbalance. Performance was measured using the macro-averaged F1 score between burned and unburned classes on the fixed test set.
We compared three types of input embeddings for the classification task. The first was TESSERA, using embeddings for 2020. The second was a task-specific composite baseline, constructed using prior knowledge of the fire event. This baseline consisted of median composites of 12 Sentinel-2 bands, with the Normalized Burn Ratio (NBR), and Normalized Difference Vegetation Index (NDVI) calculated separately for a pre-fire window (October 2019–April 2020) and a post-fire window (October 2020–April 2021), then concatenated into a single feature vector. The third type of input was Google AlphaEarth Satellite Embeddings (GSE), using publicly available annual embeddings from the GSE V1 dataset [39]. GSE and task-specific composite images were downloaded as rasters from Google Earth Engine.
A.7 Downstream Task: Above-ground biomass estimation

### v6

@@ A.2 Model Details
A.3 Pre-training Details
@@
Operationally, the process begins with the aggregation of d-pixels from all MGRS tiles into a single, comprehensive pool. A global shuffling operation is performed on this pool, a critical step to break the spatial contiguity of data from individual tiles and ensure each training batch contains a diverse mix of geographical and environmental contexts. Following this global shuffle, the data augmentation required by the Barlow Twins framework is applied.
@@
Figure 11: Data shuffling greatly improves training stability. a, Schematic of the data shuffling and loading process. D-pixels (coloured squares) from thousands of MGRS tiles are first aggregated into a global pool. A custom Rust binary performs a global shuffle on this multi-terabyte dataset before applying augmentations. The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches. b, Comparison of training loss curves. The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias. The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
A.4 Global Embedding Map Generation
@@
A.5 Scaling with data and network size
@@
A.5.1 Ablation study details
@@
Our analysis reveals several key insights into the model’s architecture and training strategy. The most critical component turns out to be the global data shuffling pipeline, which is designed to break spatial autocorrelation and present the model with geographically diverse mini-batches. Removing mixup regularization also severely degrades results. When both are removed, the model’s performance collapses, showing that the combination of global shuffling and mixup regularization is needed to prevent overfitting and to learn generalizable robust representations.
@@
A.5.2 Justification for the Number of Sampled Timesteps ($L$)
Table 5: Parameter sweep of the number of sampled timesteps ($L$). We evaluate the trade-off between F1-score on the Austrian crop classification dataset and inference speed.
$L$
Inference Speed
(samples/sec)
Encoder Param. Training Samples F1 Score (%)
15 1509 $\sim$36.4M 22M 65.04
20 1452 66.58
30 1276 71.43
40 1068 73.34
50 724 73.69
60 497 73.77
As detailed in the main paper, TESSERA is trained using sparse random temporal sampling to enforce invariance to the selection of valid observations. A critical hyperparameter in this process is $L$, the number of valid timesteps sampled. We finalized our choice at $L=40$. This decision is supported by two primary justifications:
  1. 1.
    Performance vs. Efficiency Trade-off: We conducted a comprehensive parameter sweep on the Austrian crop classification dataset to determine the optimal balance between model performance and computational cost (i.e., inference speed). As demonstrated in Table 5, setting $L=40$ achieves the best trade-off. While further increasing $L$ (e.g., to 50 or 60) results in only marginal or negligible performance gains, it increases the computational overhead during inference.
  2. 2.
    Global Data Representativeness: Our choice is also empirically grounded in global-scale data statistics. We analyzed the global distribution of valid Sentinel-2 observations acquired between 2017 and 2024 and found that the global average number of valid observations per pixel is approximately 37.6. This confirms that $L=40$ is not only efficient but also a robust and representative value for global-scale modeling, closely aligning with the average valid Sentinel-2 data availability worldwide (see Figure 12).
Figure 12: A distribution map of the number of valid Sentinel-2 observations from 2017 to 2024. Valid observations are defined as those that are free of clouds, water bodies, and snow.
Furthermore, we observed that downstream task performance is largely insensitive to the number of Sentinel-1 observations sampled (as illustrated in Figure LABEL:fig:ablation_s1_timesteps), likely due to the higher temporal density and all-weather capabilities of SAR. Therefore, for the sake of architectural consistency and to simplify the multi-modal training pipeline, we also set the number of sampled timesteps for Sentinel-1 to $L=40$.
A.6 Downstream Task Application Methodology
A.6.1 TESSERA
@@
    Download Embeddings for the Region of Interest: The GeoTessera Python library allows users to download embeddings for a desired region and year in the form of a numpy array.
@@
A.6.2 Presto & AlphaEarth
For Presto and AlphaEarth, we adopted a methodology conceptually similar to TESSERA’s, focusing on evaluating their publicly available embeddings. For Presto, we used the official Google Earth Engine (GEE) scripts provided by the authors to generate embeddings for the regions of interest. The data preprocessing was executed on GEE, and inference was subsequently performed on Google Vertex AI, loading the officially released pre-trained model weights. For AlphaEarth, we directly used the publicly available Google AlphaEarth Satellite Embedding V1 dataset [11]. In both cases, the pre-computed embeddings served as the input features. The visualization of embeddings for some downstream tasks of TESSERA, AlphaEarth, and Presto are shown in Figure 13. We then designed and trained lightweight task-specific heads (e.g., MLPs) for each downstream task, following the same principle as described for TESSERA.
Figure 13: The visualization of embeddings for some downstream tasks of TESSERA, AlphaEarth, and Presto
A.6.3 Other RSFMs
To evaluate the performance of other RSFMs, we used the standardized Pangaea evaluation framework. This framework provides a consistent protocol for benchmarking RSFMs across a suite of diverse downstream tasks. We evaluated these models on the Austrian Crop (Segmentation), PASTIS-R, Biomassters, and Borneo CHM datasets. Unlike our embedding-focused approach for TESSERA, the Pangaea protocol typically involves training task-specific decoders on top of the pre-trained encoder features. For classification tasks, a simple LinearClassifier head was attached and trained. For the regression and segmentation tasks, heavier decoders such as UperNet+LTAE or a baseline Unet were employed, as specified in the benchmark’s guidelines. The exact downstream task heads used for these models are detailed in Table 6.
Table 6: Downstream task head specification. This table lists the decoders (heads) used for different tasks across all models. Note that for each model listed, the Regression and Segmentation tasks use the same head.
Model Classification Head Regression & Segmentation Head
CROMA LinearClassifier UperNet+LTAE
DOFA LinearClassifier UperNet+LTAE
Prithvi LinearClassifier UperNet+LTAE
RemoteCLIP LinearClassifier UperNet+LTAE
SatlasNet LinearClassifier UperNet+LTAE
Scale-MAE LinearClassifier UperNet+LTAE
SpectralGPT LinearClassifier UperNet+LTAE
Galileo LinearClassifier UperNet+LTAE
Skysense LinearClassifier UperNet+LTAE
ViT Baseline LinearClassifier UperNet+LTAE
UNet Baseline LinearClassifier Unet
Presto LinearClassifier Unet
AlphaEarth LinearClassifier Unet
TESSERA LinearClassifier Unet
A.7 Downstream Task: Austrian Crop Classification
@@
We used the publicly available INVEKOS dataset for Austria, focusing on the 2022 growing season [1]. The data set originally contained 154 crop types, which we grouped into 17 broader classes (for example, combining sugar beet and feed beet) based on phenological similarity and sample availability to ensure robust training and evaluation.
When constructing the pixel-level classification dataset, we use fields as the unit for data splitting. Each field is assigned a unique identifier along with its corresponding area. We allocate pixels from fields constituting X% of the total area to the training set, while the remaining (1-X)% of the total field area is divided such that 1/7 is used for validation and the rest for testing. For training, validation, and test splits, we ensure that all 17 crop types are represented in each split.
For the patch-based dataset designed for segmentation tasks, we first divide the region of interest vertically into five equal parts. From the topmost section, we extract 1000 patches of 640m × 640m as the test set. An equal number of patches with the same dimensions are cropped from the section immediately below for the validation set. From the remaining three sections, 3000 patches of the same size are sampled to form the training set. All patches maintain a spatial resolution of 10 meters.
@@
  • AlphaEarth + MLP: To include another key foundation model, we used the publicly available Google AphaEarth Satellite Embedding V1 dataset [11]. These embeddings were similarly fed into an MLP head identical to the one used for TESSERA.
  • Random Forest: As a traditional baseline, we trained a Random Forest (RF) classifier directly on raw time-series data. For each pixel, we used all available Sentinel-1 (2 polarizations) and Sentinel-2 (10 spectral bands) observations throughout the year. The temporal and spectral/polarization dimensions were flattened and concatenated to form a single 1256-dimensional feature vector. The RF model was configured with 100 trees, while all other hyperparameters retained their default values as implemented in the scikit-learn library.
@@
  • For the foundation models capable of generating pixel-wise embeddings, namely TESSERA, Presto [88], and AlphaEarth [11], a distinct workflow was used. We first construct patches from their pixel-level embeddings (e.g., a data cube of size $H\times W\times C$). These patches were then fed into a standard UNet architecture [73], which learns to model the spatial relationships between the embeddings to produce a segmentation map. It is noteworthy that while TESSERA and Presto’s base embeddings do not explicitly model spatial context, AlphaEarth’s do. However, their shared ability to generate pixel-level outputs makes them suitable for this UNet-based downstream approach.
@@
Figure 14: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation. Each row corresponds to a different patch size (32x32, 64x64, 128x128). The first column shows the ground truth segmentation. For each model (TESSERA, Presto, and AlphaEarth), two columns are displayed: a PCA visualization of the input embeddings for the patch (labelled ’Embedding’) and the final predicted segmentation map (labelled ’Predicted’).
@@
Figure 15: UMAP visualization of the embedding spaces for TESSERA, Presto, and AlphaEarth for 17 crop classes. TESSERA’s embeddings exhibit clearer separation and more coherent clustering visually.
The 2D visualizations shown in Figure 15 were generated by applying the Uniform Manifold Approximation and Projection (UMAP) [60] algorithm to the embeddings of every pixel within the Austrian study area. UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology. It constructs a weighted k-nearest neighbour graph in the original high-dimensional space, then optimizes a low-dimensional projection to primarily preserve local distances and local structure. UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space. In this work, the entire embedding map (for example, a shape array $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for the TESSERA, Presto, and AlphaEarth embeddings.
@@
Impact of Seasonal Windows
We analyze the impact of using different temporal windows on the model’s performance for the downstream task of crop classification on the Austrian dataset. As shown in Table 8, we conduct an ablation by systematically excluding one of the four meteorological seasons (Spring, Summer, Autumn, Winter) from the input time series. The results reveal that, for this specific regional dataset, excluding time steps from the winter season improves the classification F1 score from 78.94% (all seasons included) to $\mathbf{79.83\%}$. This suggests that the dormant winter period, which is typically characterized by low vegetation phenological activity and is prone to cloud cover, snow, or reduced solar zenith angle effects, introduces a substantial amount of non-discriminative noise. By excluding the winter data, the model can focus its attention and resources on the most informative growing seasons, which allows for the learning of a more concise and effective temporal representation for crop type identification.
Table 8: Ablation on the Austrian Crop dataset: Seasonal Data Inclusion. We report the Validation F1 score (higher $\uparrow$ is better) for the crop classification task. The columns for Spring to Winter use  to indicate inclusion and  to indicate exclusion. Parentheses show change vs. the Full Model.
Configuration Spring Summer Autumn Winter F1 ($\uparrow$)
Full Model (Baseline) 78.94
w/o Spring 78.62 (-0.32)
w/o Summer 78.10 (-0.84)
w/o Autumn 78.22 (-0.72)
w/o Winter 79.83 (+0.89)
A.8 Downstream Task: Canopy height estimation
@@
TESSERA, Presto [88], and AlphaEarth [11] embeddings were compared by training identical models using each representation as input, with all other settings held constant. Each model was trained for 200 epochs on the same data split, and the version achieving the best validation performance was selected for final evaluation on the held-out test set.
To benchmark performance against classical remote sensing approaches applied at the global scale, we compared the predictions against a global canopy height product: Tolan et al. [2024][87] (Meta). These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS. Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region, which is reasonable given that the area was not logged or burned.
Figure 16: Comparison of canopy height estimation performance across four models. Predicted canopy height (y-axis) is compared against airborne LiDAR reference measurements (x-axis) in Danum Valley, Borneo. The figure shows scatter-density plots for four model variants trained using different input representations: TESSERA, Presto, AlphaEarth, and the DINOv2 Global CHM product (renamed from the original Meta Global CHM). All models were evaluated using the same LiDAR reference dataset for consistency. TESSERA achieves the highest coefficient of determination and lowest RMSE, demonstrating superior performance in structurally complex tropical canopies. In contrast, the DINOv2 Global CHM product exhibits strong underestimation, large negative bias, and a near-zero or negative correlation, reflecting substantial saturation at lower canopy heights. A 1:1 reference line (blue) is included in each subplot for comparison.
@@
A.9 Downstream Task: Above-ground biomass estimation